Carla's company makes water heaters. These heaters last an average of 60 months, with a standard deviation of 8 months, before needing to be serviced. Her accountants estimate that the company can afford to replace 10% of all water heaters and still make a tidy profit.

Rounding to the nearest whole number, how many months should the warranty be?
Answer #1

Given that,

mean = = 60

standard deviation = =8

Using standard normal table,

P(Z < z) = 10%

= P(Z < z) = 0.10  

= P(Z <-1.28) = 0. 10

z = -1.28

Using z-score formula  

x= z * +

x= -1.28 *8+60

x= 49.76

x=50